a farmer borrow $3800 for 9 months from a bank. The discounted the loan at 7%
I know that interest is $199.50 and amount of money the farmer recieved is $3600.50, but can't figure out the actual interest rate which is 7.39%

Answer by mananth(16946)

I= p*n*r
I= 3800*3/4 * r
199.5 = (3800-199.5) * 3/4*r
r= (199.5*4)/(3*3600.50)
r=0.07387
r= 7.387% round off